Suppose a large shipment of compact discs contained 19% defectives.

If a sample of size 303 is selected, what is the probability that the sample proportion will be less than 17%? Round your answer to four decimal places.
Answer #1

P = 0.19

n = 303

np = 57.57 > 10

n ( 1 - P ) = 245.43 > 10

Sampling distribution of is approximate to normal

=  

X ~ N ( µ = 0.19 , σ = 0.0225 )
P ( X < 0.17 )
Standardizing the value
Z = ( X - µ ) / σ
Z = ( 0.17 - 0.19 ) / 0.0225
Z = -0.8889
P ( ( X - µ ) / σ ) < ( 0.17 - 0.19 ) / 0.0225 )
P ( X < 0.17 ) = P ( Z < -0.8889 )
P ( X < 0.17 ) = 0.1870

P ( < 0.17 ) = 0.1870